OpenStudy (shaeelynn):

The sum Sn of finite arithmetic series of n terms is Sn= n/2(A1+An) where A1 is the first term and An is the nth term. a) show that Sn=na1+ n(n-1)/2 d by replacing An with its value in term of A1, n, and d in the above formula b) explain why Sn=na1+ n(n-1) d makes sense by explaining how you can extract each of na1 and n(n-1)/2 d from the sum A1+(A1+d)+(A1+2d)+...+(a1+(n-1)d)

OpenStudy (xguardians):

@rebeccaxhawaii , practice for u bb

rebeccaxhawaii (rebeccaxhawaii):

can you take a screenshot of the question.

OpenStudy (shaeelynn):

someone please help i have no idea how to do this

jimthompson5910 (jim_thompson5910):

Hint: \[\Large S_n = \frac{n}{2}\left(a_1 + a_n\right)\] \[\Large S_n = \frac{n}{2}\left(a_1 + {\color{red}{a_n}}\right)\] \[\Large S_n = \frac{n}{2}\left(a_1 + {\color{red}{a_1+d(n-1)}}\right)\] I'm using the nth term formula (for arithmetic sequences) which is \[\Large a_n = a_1+d(n-1)\]

OpenStudy (shaeelynn):

it still makes no sense at all to me

jimthompson5910 (jim_thompson5910):

Do you see how I replaced the \(\Large a_n\) with \(\Large a_1+d(n-1)\) ?

OpenStudy (dessyj1):

The first question is trying to get us to show that Sn=n(a1+an)/2=na1+n(n-1)d/2.

OpenStudy (dessyj1):

do you agree?

OpenStudy (shaeelynn):

yes i do and yes i agree that it is showing that

OpenStudy (dessyj1):

If you expand the sum Sn - the sum up to the nth term - you would get some thing that looks like this:

OpenStudy (dessyj1):

Finally, you just need to substitute an into Sn=n(a1+an)/2 and it should produce Sn= na1+n(n-1)d/2 if you expand the terms.